# Tilting representations of finite groups of Lie type

### v1

@@ 2.4. Action of $\mathbf{T}^{\vee}$
Recall that monodromic sheaves are sheaves that are locally constant along $\mathbf{T}\times\mathbf{T}$-orbits and tame local systems, we refer to [Ete24, Section 2] for a discussion. Since all strata are classifying spaces the monodromic part is clear. Monodromic sheaves are equipped with a canonical action of $\pi_{1}^{t}(\mathbf{T}\times\mathbf{T})$, the tame fundamental group of $\mathbf{T}\times\mathbf{T}$. This groups is isomorphic to $X_{*}(\mathbf{T}\times\mathbf{T})\otimes\hat{\mathbb{Z}}^{(p)}(1)$. After fixing a trivialization of the roots of $1$ in $\overline{\mathbb{F}}_{q}$ we get a topological generator of ^$\hat{\mathbb{Z}}^{(p)}(1)$ and an action of $\Lambda[X_{*}(\mathbf{T}\times\mathbf{T})]=\mathcal{O}(\mathbf{T}^{\vee})\otimes_{\Lambda}\mathcal{O}(\mathbf{T}^{\vee})$ on monodromic sheaves. We first restrict this action to an action of $\mathcal{O}(\mathbf{T}^{\vee})^{\mathbf{W}}\otimes_{\Lambda}\mathcal{O}(\mathbf{T}^{\vee})^{\mathbf{W}}$. Finally we observe that the $\mathrm{Ad}_{\mathrm{F}}(\mathbf{T})$-equivariance shows that the map
@@
factors through the quotient of $\mathcal{O}({\mathbf{T}}^{\vee})^{\mathbf{W}}\otimes_{\Lambda}\mathcal{O}({\mathbf{T}}^{\vee})^{\mathbf{W}}$ by the ideal $(f-\mathrm{F}^{\vee}(f),f\in\mathcal{O}(\mathbf{T}^{\vee})^{\mathbf{W}})$, that is, by $\mathcal{O}(\mathbf{T}^{\vee}\sslash\mathbf{W})^{\mathrm{F}^{\vee}})$. ∎
@@
2.5. Mixed Deligne-Lusztig category $\mathcal{O}$
We equip the stack $\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}$ with the endomorphism $\mathrm{F}^{\delta}$ where $\delta$ is minimal such that the pair $(\mathbf{G},\mathrm{F}^{\delta})$ is a split reductive group over $\mathbb{F}_{q^{\delta}}$. All mixed structures are done with respect to this Frobenius. Furthermore, the Bruhat stratification is stable under $\mathrm{F}^{\delta}$. We define the category $\mathcal{O}^{\mathrm{DL},\mathrm{mix}}$ to be the full subcategory of constructible mixed $\overline{\mathbb{Q}}_{\ell}$-sheaves on $\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}$ generated as a triangulated category by Tate twists of pure $\mathrm{IC}$-sheaves of weight 00. There is a natural forgetful functor
@@ 2.5. Mixed Deligne-Lusztig category $\mathcal{O}$
$[-]:\mathcal{H}\simeq K_{0}(\mathcal{O}^{\mathrm{DL},\mathrm{mix},\mathrm{unip}}_{\overline{\mathbb{Q}}_{\ell}}).$
@@
$[\Delta_{w,1}^{\mathrm{mix}}]=H_{w},[\nabla_{w,1}^{\mathrm{mix}}]=H_{w^{-1}}^{-1},[T_{w,1}^{\mathrm{mix}}]=\underline{H}_{w},[\mathrm{IC}_{w,1}^{\mathrm{mix}}]=\underline{\tilde{H}}_{w},[-(\frac{1}{2})]=v^{-1}[-].$
@@
satisfying $[\Delta_{w,1}^{\mathrm{mix}}]=H_{w}$. Similarly, there are isomorphisms between $\mathcal{H}$ and the Grothendieck groups of the two mixed realizations of the Hecke categories $\mathcal{H}^{\mathrm{eq},\mathrm{unip},\mathrm{mix}}$ and $\mathcal{H}^{\mathrm{mon},\mathrm{unip},\mathrm{mix}}$ constructed in [BY13], that is we have
@@
Since the pushforward $\mathcal{H}^{\mathrm{mon},\mathrm{unip},\mathrm{mix}}\to\mathcal{O}^{\mathrm{DL},\mathrm{unip},\mathrm{mix}}$ sends standard (resp. costandard, resp. tilting) to standard (resp. costandard, resp. tilting), we deduce from the categorification theorem of loc. cit. that $[\nabla_{w,1}^{\mathrm{mix}}]=H_{w^{-1}}^{-1}$ and $[T_{w,1}^{\mathrm{mix}}]=\underline{H}_{w}$. Similarly, the pullback $\mathcal{H}^{\mathrm{eq},\mathrm{unip},\mathrm{mix}}\to\mathcal{O}^{\mathrm{DL},\mathrm{unip},\mathrm{mix}}$ sends standard (resp. costandard, resp. simple) to standard (resp. costandard, resp. simple) which allows us to deduce that $[\mathrm{IC}_{w,1}^{\mathrm{mix}}]=\underline{\tilde{H}}_{w}$. ∎
@@
The reader should note that while there is a ring structure on $\mathcal{H}_{q}$, there is a priori no ring or action defined on the RHS. However this equivalence is linear over $\mathbb{Z}[v^{\pm 1}]$.
@@
()¯$\overline{(-)}:\mathcal{H}_{\mathrm{mon}}\to\mathcal{H}_{\mathrm{mon}},v^{n}H_{w}1_{\chi}\mapsto v^{-n}H^{-1}_{w^{-1}}1_{\chi}.$
@@
$[\Delta_{w,\chi}^{\mathrm{mix}}]=H_{w,\chi},[\nabla_{w,\chi}^{\mathrm{mix}}]=H_{w^{-1},\chi}^{-1},[\mathrm{IC}_{w,\chi}^{\mathrm{mix}}]=\underline{\tilde{H}}_{w,\chi}.$
@@
$[T_{w,\chi}^{\mathrm{mix}}]=\underline{H}_{w,\chi}.$
The proof of this conjecture in the unipotent case relied on the study of free monodromic mixed unipotent sheaves as done in [BY13]. We did not find a source in the literature that does this in the non unipotent mixed case. In the thesis of Gouttard [Gou22], the author studies the non-unipotent case but not in the mixed setting, it can be deduced from the categorification theorems of loc. cit. that
$[T_{w,\chi}](1)=\underline{H}_{w,\chi}(1).$
@@ 2.6. Relation with category $\mathcal{O}$
We consider the categories of unipotent free monodromic sheaves on $\mathbf{U}\backslash\mathbf{G}/\mathbf{U}$ as constructed in [BY13] and [Ete24] for $\Lambda=\overline{\mathbb{Q}}_{\ell}$ and $\Lambda=\overline{\mathbb{Z}}_{\ell}$, we denote these categories by ¯$\mathcal{H}^{\mathrm{mon},\mathrm{unip}}_{\overline{\mathbb{Z}}_{\ell}}$ and ¯$\mathcal{H}^{\mathrm{mon},\mathrm{unip}}_{\overline{\mathbb{Q}}_{\ell}}$. Moreover, there is an inversion of $\ell$-functor
@@
Let $T\in\mathcal{H}^{\mathrm{mon},\mathrm{unip},\overline{\mathbb{Z}}_{\ell}}$ be a free monodromic tilting object as constructed in [Ete23] and [Gou22], by [Ete24] the object $\mathrm{inv}_{\ell}T\in\mathcal{H}^{\mathrm{mon},\mathrm{unip},\overline{\mathbb{Q}}_{\ell}}$ also free monodromic tilting. Furthermore the following relations hold by definition of the functor $\mathrm{inv}_{\ell}$. We have
@@
The decomposition numbers that appear in Lemma 2.6.1 control the difference between the usual Kazhdan-Lusztig basis in $\mathcal{H}$ and the $\ell$-Kazhdan-Lusztig basis (which depends on $\mathbf{G}$ and not just on $\mathbf{W}$), firstly defined in the case of crystallographic Coxeter groups in [JMW14]. We refer to [JW17] for an account on this basis.
@@
For a fixed $\mathbf{G}$, there exists $\ell_{0}$ large enough such that for all $\ell\geq\ell_{0}$ the Kazhdan-Lusztig and $\ell$-Kazhdan-Lusztig basis coincide. For such $\ell$, we have $n_{v,w,1}=0$ for all $v<w$.
@@
Predicting the behaviour of this basis and in particular when it differs from the usual Kazhdan-Lusztig basis is a very hard problem. In [GJW23], the authors construct an algorithm to compute this basis.
@@ 3.1. Horocycle correspondence
[…]
@@
    There is an isomorphism of functors $\mathrm{ch}j_{w,!}=R_{w}[\ell(w)]$ where the RHS is the Deligne-Lusztig induction functor.
@@
The first point is done in the original paper of Deligne and Lusztig [DL76] if $\Lambda=\overline{\mathbb{Q}}_{\ell}$. The integral and the modular cases are done in [BM89, Theorem 2.2]. In view of Theorem 3.1.1, the second point follows from the definition of the Lusztig series and the compatibility between the Deligne-Lusztig functors and the functors $\mathrm{ch}$ and $\mathrm{hc}$ of theorem 3.1.1. ∎
@@ 3.2. Tilting representations
    Assume that all the Deligne-Lusztig varieties of $(\mathbf{G},\mathrm{F})$ are affine, then $\mathrm{ch}(T)$ is concentrated in degree 00 and is a projective object of $\mathrm{Rep}_{\Lambda}\mathbf{G}^{\mathrm{F}}$.
@@
Since the map $\alpha$ is proper and smooth and the map $\beta$ is smooth the functor $\mathrm{ch}$ has a continuous right adjoint hence $\mathrm{ch}$ preserves compact objects. Since tilting sheaves are compact by Proposition 2.1.3, the objects $\mathrm{ch}(T)$ are compact. Assume now that the Deligne-Lusztig varieties are affine. By Theorem 3.1.1, it follows that $\mathrm{ch}(\Delta)\in\mathrm{D}^{\geq 0}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}})$ (resp. $\mathrm{ch}(\nabla)\in\mathrm{D}^{\leq 0}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}})$). Since $T$ has both a $\Delta$ and a $\nabla$-filtration, we get that $\mathrm{ch}(T)\in\mathrm{D}^{\heartsuit}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}})$. Hence $\mathrm{ch}(T)$ is a representation of $\mathbf{G}^{\mathrm{F}}$ concentrated in degree 00, which is compact in $\mathrm{D}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}})$. Since $\Lambda[\mathbf{G}^{\mathrm{F}}]$ is a symmetric algebra, it follows that the representation $\mathrm{ch}(T)$ is projective. ∎
Remark 3.2.3.
The affineness of the Deligne-Lusztig varieties is a very mild hypothesis, indeed, if $G$ is defined over $\mathbb{F}_{q}$, then this hypothesis is satisfied if $q>h$ by [DL76, Theorem 9.7] where $h$ is the Coxeter number of $G$.
@@ 4.1. The Alvis–Curtis dual of an $\mathrm{IC}$ representation
We consider the Alvis-Curtis duality functor, see [DL82]
@@
[…]
@@
Tracing the definitions, the statement of [Lus84, Theorem 3.8] is precisely this equality evaluated at $\underline{H}_{w}$ since these elements form a basis of $\mathcal{H}$ and our formulation of the theorem follows. ∎
@@
$d(\mathrm{ch}([\mathrm{IC}_{w,1}^{\mathrm{mix}}]))=\pm\mathrm{ch}(T_{w,1}^{\mathrm{mix}}).$
@@
We have restricted ourselves to the unipotent case as this will be enough to prove the conjecture of Dudas-Malle that we are interested in. However it should be noted that, assuming Conjecture 2.5.9, a non-unipotent statement as Corollary 4.1.4 could be deduced. Indeed, Theorems 4.1.2 and 4.1.3 are both shown in [Lus84] in the non-unipotent case.

### v2

@@ 2.4. Action of $\mathbf{T}^{\vee}$
Recall that monodromic sheaves are sheaves that are locally constant along $\mathbf{T}\times\mathbf{T}$-orbits and tame local systems, we refer to [Ete24, Section 2] for a discussion. Since all strata are classifying spaces the monodromic part is clear. Monodromic sheaves are equipped with a canonical action of $\pi_{1}^{t}(\mathbf{T}\times\mathbf{T})$, the tame fundamental group of $\mathbf{T}\times\mathbf{T}$. This groups is isomorphic to $X_{*}(\mathbf{T}\times\mathbf{T})\otimes\hat{\mathbb{Z}}^{(p)}(1)$. After fixing a trivialization of the roots of $1$ in $\overline{\mathbb{F}}_{q}$ we get a topological generator of $\hat{\mathbb{Z}}^{(p)}(1)$ and an action of $\Lambda[X_{*}(\mathbf{T}\times\mathbf{T})]=\mathcal{O}(\mathbf{T}^{\vee})\otimes_{\Lambda}\mathcal{O}(\mathbf{T}^{\vee})$ on monodromic sheaves. We first restrict this action to an action of $\mathcal{O}(\mathbf{T}^{\vee})^{\mathbf{W}}\otimes_{\Lambda}\mathcal{O}(\mathbf{T}^{\vee})^{\mathbf{W}}$. Finally we observe that the $\mathrm{Ad}_{\mathrm{F}}(\mathbf{T})$-equivariance shows that the map
@@
factors through the quotient of $\mathcal{O}({\mathbf{T}}^{\vee})^{\mathbf{W}}\otimes_{\Lambda}\mathcal{O}({\mathbf{T}}^{\vee})^{\mathbf{W}}$ by the ideal $(f-\mathrm{F}^{\vee}(f),f\in\mathcal{O}((\mathbf{T}^{\vee})^{\mathbf{W}})$, that is, by $\mathcal{O}(\mathbf{T}^{\vee}\sslash\mathbf{W})^{\mathrm{F}^{\vee}})$. ∎
@@
2.5. Mixed Deligne–Lusztig category $\mathcal{O}$
We equip the stack $\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}$ with the endomorphism $\mathrm{F}^{\delta}$ where $\delta$ is minimal such that the pair $(\mathbf{G},\mathrm{F}^{\delta})$ is a split reductive group over $\mathbb{F}_{q^{\delta}}$. All mixed structures are done with respect to this Frobenius. Furthermore, the Bruhat stratification is stable under $\mathrm{F}^{\delta}$. We define the category $\mathcal{O}^{\mathrm{DL},\mathrm{mix}}$ to be the full subcategory of constructible mixed $\overline{\mathbb{Q}}_{\ell}$-sheaves on $\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}$ generated as a triangulated category by Tate twists of pure $\mathrm{IC}$-sheaves of weight 0. There is a natural forgetful functor
@@ 2.5. Mixed Deligne–Lusztig category $\mathcal{O}$
$\mathcal{H}\simeq K_{0}(\mathcal{O}^{\mathrm{DL},\mathrm{mix},\mathrm{unip}}_{\overline{\mathbb{Q}}_{\ell}}).$
@@
$\Delta_{w,1}^{\mathrm{mix}}]=H_{w},[\nabla_{w,1}^{\mathrm{mix}}]=H_{w^{-1}}^{-1},[T_{w,1}^{\mathrm{mix}}]=\underline{H}_{w},[\mathrm{IC}_{w,1}^{\mathrm{mix}}]=\underline{\tilde{H}}_{w},[-(\frac{1}{2})]=v^{-1}[-].$
@@
satisfying $\Delta_{w,1}^{\mathrm{mix}}]=H_{w}$. Similarly, there are isomorphisms between $\mathcal{H}$ and the Grothendieck groups of the two mixed realizations of the Hecke categories $\mathcal{H}^{\mathrm{eq},\mathrm{unip},\mathrm{mix}}$ and $\mathcal{H}^{\mathrm{mon},\mathrm{unip},\mathrm{mix}}$ constructed in [BY13], that is we have
@@
Since the pushforward $\mathcal{H}^{\mathrm{mon},\mathrm{unip},\mathrm{mix}}\to\mathcal{O}^{\mathrm{DL},\mathrm{unip},\mathrm{mix}}$ sends standard (resp. costandard, resp. tilting) to standard (resp. costandard, resp. tilting), we deduce from the categorification theorem of loc. cit. that $\nabla_{w,1}^{\mathrm{mix}}]=H_{w^{-1}}^{-1}$ and $T_{w,1}^{\mathrm{mix}}]=\underline{H}_{w}$. Similarly, the pullback $\mathcal{H}^{\mathrm{eq},\mathrm{unip},\mathrm{mix}}\to\mathcal{O}^{\mathrm{DL},\mathrm{unip},\mathrm{mix}}$ sends standard (resp. costandard, resp. simple) to standard (resp. costandard, resp. simple) which allows us to deduce that $\mathrm{IC}_{w,1}^{\mathrm{mix}}]=\underline{\tilde{H}}_{w}$. ∎
@@
The reader should note that while there is a ring structure on $\mathcal{H}_{q}$, there is a priori no ring structure or action defined on the RHS. However this equivalence is linear over $\mathbb{Z}[v^{\pm 1}]$.
@@
$\overline{(-)}:\mathcal{H}_{\mathrm{mon}}\to\mathcal{H}_{\mathrm{mon}},v^{n}H_{w}1_{\chi}\mapsto v^{-n}H^{-1}_{w^{-1}}1_{\chi}.$
@@
$\Delta_{w,\chi}^{\mathrm{mix}}]=H_{w,\chi},[\nabla_{w,\chi}^{\mathrm{mix}}]=H_{w^{-1},\chi}^{-1},[\mathrm{IC}_{w,\chi}^{\mathrm{mix}}]=\underline{\tilde{H}}_{w,\chi}.$
@@
$T_{w,\chi}^{\mathrm{mix}}]=\underline{H}_{w,\chi}.$
The proof of this conjecture in the unipotent case relied on the study of free monodromic mixed unipotent sheaves as done in [BY13]. We did not find a source in the literature that does this in the non-unipotent mixed case. In the thesis of Gouttard [Gou22], the author studies the non-unipotent case but not in the mixed setting, it can be deduced from the categorification theorems of loc. cit. that
$T_{w,\chi}](1)=\underline{H}_{w,\chi}(1).$
@@ 2.6. Relation with category $\mathcal{O}$
We consider the categories of unipotent free monodromic sheaves on $\mathbf{U}\backslash\mathbf{G}/\mathbf{U}$ as constructed in [BY13] and [Gou22] for $\Lambda=\overline{\mathbb{Q}}_{\ell}$ and $\Lambda=\overline{\mathbb{Z}}_{\ell}$, we denote these categories by $\mathcal{H}^{\mathrm{mon},\mathrm{unip}}_{\overline{\mathbb{Z}}_{\ell}}$ and $\mathcal{H}^{\mathrm{mon},\mathrm{unip}}_{\overline{\mathbb{Q}}_{\ell}}$. Moreover, there is an inversion of $\ell$-functor
@@
Let $T\in\mathcal{H}^{\mathrm{mon},\mathrm{unip},\overline{\mathbb{Z}}_{\ell}}$ be a free monodromic tilting object, then the object $\mathrm{inv}_{\ell}T\in\mathcal{H}^{\mathrm{mon},\mathrm{unip},\overline{\mathbb{Q}}_{\ell}}$ also free monodromic tilting. Furthermore the following relations hold by definition of the functor $\mathrm{inv}_{\ell}$. We have
@@
The decomposition numbers that appear in Lemma 2.6.1 control the difference between the usual Kazhdan–Lusztig basis in $\mathcal{H}$ and the $\ell$-Kazhdan–Lusztig basis (which depends on $\mathbf{G}$ and not just on $\mathbf{W}$), firstly defined in the case of crystallographic Coxeter groups in [JMW14]. We refer to [JW17] for an account on this basis.
@@
For a fixed $\mathbf{G}$, there exists $\ell_{0}$ large enough such that for all $\ell\geq\ell_{0}$ the Kazhdan-Lusztig and $\ell$-Kazhdan–Lusztig basis coincide. For such $\ell$, we have $n_{v,w,1}=0$ for all $v<w$.
@@
Predicting the behaviour of this basis and in particular when it differs from the usual Kazhdan–Lusztig basis is a very hard problem. In [GJW23], the authors construct an algorithm to compute this basis.
@@ 3.1. Horocycle correspondence
$\frac{\mathbf{G}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{G}}\xleftarrow{\alpha}\frac{\mathbf{G}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{B}}\xrightarrow{\beta}\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}.$
@@
    There is an isomorphism of functors $\mathrm{ch}j_{w,!}=R_{w}[\ell(w)]$ where the RHS is the Deligne–Lusztig induction functor.
@@
The first point is done in the original paper of Deligne and Lusztig [DL76] if $\Lambda=\overline{\mathbb{Q}}_{\ell}$. The integral and the modular cases are done in [BM89, Theorem 2.2]. In view of Theorem 3.1.1, the second point follows from the definition of the Lusztig series and the compatibility between the Deligne–Lusztig functors and the functors $\mathrm{ch}$ and $\mathrm{hc}$ of theorem 3.1.1. ∎
@@ 3.2. Tilting representations
Lemma 3.2.2.
All the complexes $\mathrm{ch}(\bigoplus_{\chi}\Delta_{w,\chi})$ are concentrated in nonnegative degree.
Proof.
This is a consequence of Artin vanishing as all Deligne–Lusztig varieties are affine. If $q$ is large enough then this is [DL76, Theorem 9.7], the general case is done in [BHL25, Corollary 3.12]. ∎
Remark 3.2.3.
Another proof of Lemma 3.2.2 is done in [Zhu25] and is independent of [BHL25].
@@
    The complex $\mathrm{ch}(T)$ is concentrated in degree 0 and is a projective object of $\mathrm{Rep}_{\Lambda}\mathbf{G}^{\mathrm{F}}$.
@@
Since the map $\alpha$ is proper and smooth and the map $\beta$ is smooth the functor $\mathrm{ch}$ has a continuous right adjoint hence $\mathrm{ch}$ preserves compact objects. Since tilting sheaves are compact by Proposition 2.1.3, the objects $\mathrm{ch}(T)$ are compact.
By Theorem 3.1.1 and Lemma 3.2.2, it follows that $\mathrm{ch}(\Delta)\in\mathrm{D}^{\geq 0}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}})$ (resp. $\mathrm{ch}(\nabla)\in\mathrm{D}^{\leq 0}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}})$). Since $T$ has both a $\Delta$ and a $\nabla$-filtration, we get that $\mathrm{ch}(T)\in\mathrm{D}^{\heartsuit}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}})$. Hence $\mathrm{ch}(T)$ is a representation of $\mathbf{G}^{\mathrm{F}}$ concentrated in degree 0, which is compact in $\mathrm{D}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}})$. Since $\Lambda[\mathbf{G}^{\mathrm{F}}]$ is a symmetric algebra, it follows that the representation $\mathrm{ch}(T)$ is projective. ∎
@@ 4.1. The Alvis–Curtis dual of an $\mathrm{IC}$ representation
We consider the Alvis–Curtis duality functor, see [DL82]
@@
    $\mathcal{H}=K_{0}(\mathcal{O}^{\mathrm{DL},\mathrm{mix},\mathrm{unip}}_{\overline{\mathbb{Q}}_{\ell}})\xrightarrow{\mathrm{ch}}K_{0}(\mathrm{Rep}^{\mathrm{mix}}_{\overline{\mathbb{Q}}_{\ell}}\mathbf{G}^{\mathrm{F}})\xrightarrow{\gamma}\mathbb{Q}(v)\otimes R(\mathbf{G}^{\mathrm{F}})$, we shall denote this map by $\mathrm{ch}$,
@@
Tracing the definitions, the statement of [Lus84, Theorem 3.8] is precisely this equality evaluated at $\underline{H}_{w}$. Since these elements form a basis of $\mathcal{H}$, the theorem follows. ∎
@@
$d(\mathrm{ch}([\mathrm{IC}_{w,1}^{\mathrm{mix}}]))=\pm\mathrm{ch}([T_{w,1}^{\mathrm{mix}}]).$
@@
We have restricted ourselves to the unipotent case, as this suffices to prove the conjecture of Dudas–Malle. However it should be noted that, assuming Conjecture 2.5.9, a non-unipotent variant of Corollary 4.1.4 could be deduced. Indeed, Theorems 4.1.2 and 4.1.3 are both shown in [Lus84] in the non-unipotent case.
4.2. Eigenvalues of Frobenius
We make a quick reminder on the theory of eigenvalues of Frobenius.
Theorem 4.2.1 ([Lus84, Theorem 3.8.1]).
Let $\rho\in\mathrm{Irr}_{\overline{\mathbb{Q}}_{\ell}}^{\mathrm{unip}}\mathbf{G}^{\mathrm{F}}$ be an irreducible unipotent representation. There exists a root of unity $\zeta_{\rho}\in\overline{\mathbb{Q}}_{\ell}^{\times}$ satisfying the following property, for all $A\in\mathcal{O}^{\mathrm{DL},\mathrm{unip},\mathrm{mix}}$, the eigenvalues of $\mathrm{F}^{\delta}$ on the $\rho$-isotypic component of $\mathrm{ch}(A)$, are of the form $q^{\frac{i}{2}}\zeta$ for $i\in\mathbb{Z}$.
Proof.
This exact statement is proved in [Lus84, Theorem 3.8.1] when $A=\Delta_{w,1}^{\mathrm{mix}}$ (and thus $\mathrm{ch}(A)=\mathrm{R\Gamma}_{c}(X(w),\overline{\mathbb{Q}}_{\ell})[\ell(w)]$). Since these objects generate the category $\mathcal{O}^{\mathrm{DL},\mathrm{unip},\mathrm{mix}}$, the statement follows. ∎
Let $\rho$ be as in Theorem 4.2.1, and let $d(\rho)\in\mathrm{Irr}_{\overline{\mathbb{Q}}_{\ell}}^{\mathrm{unip}}\mathbf{G}^{\mathrm{F}}$ be the Alvis–Curtis dual of $\rho$ (written without signs).
Lemma 4.2.2.
The following equality holds
$\zeta_{d(\rho)}=\zeta_{\rho}.$
Proof.
By [Lus78, 3.33] (see also [GM03, Lemma 4.2]), the map
$\rho\mapsto\zeta_{\rho}$
is constant on Harish-Chandra series. Since Harish-Chandra series are invariant under Alvis–Curtis duality, the statement follows. ∎
Recall (see 1.3) that we denote by $\mathrm{D}_{c}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}},\overline{\mathbb{Q}}_{\ell})^{\mathrm{Weil}}$ the category of pairs $(\rho,\alpha)$ where $\rho\in\mathrm{D}_{c}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}},\overline{\mathbb{Q}}_{\ell})$ and $\alpha$ is an isomorphism $\mathrm{F}^{\delta,*}\rho\simeq\rho$. For $\lambda\in\overline{\mathbb{Q}}_{\ell}$ and $(\rho,\alpha)\in\mathrm{D}_{c}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}},\overline{\mathbb{Q}}_{\ell})^{\mathrm{Weil}}$, we denote by
$\rho[\lambda]$
$\rho[\overline{\lambda}]=\bigoplus_{\lambda}\rho[\lambda]$
where the direct sum is indexed by all $\lambda\in\overline{\mathbb{Z}}_{\ell}$ which reduce to $\overline{\lambda}\in\overline{\mathbb{F}}_{\ell}$.
Lemma 4.2.3.
Let $A\in\mathcal{O}^{\mathrm{DL},\mathrm{unip},\mathrm{mix}}$, then there is an equality in $K_{0}(\mathrm{Rep}_{\overline{\mathbb{Q}}_{\ell}}(\mathbf{G}^{\mathrm{F}}))$ for all $\lambda\in\overline{\mathbb{Q}}_{\ell}^{\times}$,
$d(\mathrm{ch}(A)[\lambda])=(d(\mathrm{ch}(A))[\lambda].$
Proof.
It is enough to prove the statement for a generating collection of sheaves in $\mathcal{O}^{\mathrm{DL},\mathrm{unip},\mathrm{mix}}$, we will prove it for $\mathrm{IC}_{w}^{\mathrm{mix}}$. By Theorem 4.2.1, we can assume that $\lambda=q^{\frac{i}{2}}\zeta$ for some root of unity $\zeta$. Since the equality of Corollary 4.1.4 is $\mathbb{Q}(v)$-linear, we have
$\sum_{\zeta^{\prime}}d(\mathrm{IC}_{w}^{\mathrm{mix}}[q^{\frac{i}{2}}\zeta^{\prime}])=\sum_{\zeta^{\prime}}d(\mathrm{IC}_{w}^{\mathrm{mix}})[q^{\frac{i}{2}}\zeta^{\prime}].$ (1)
Let $\rho\in\mathrm{Irr}_{\overline{\mathbb{Q}}_{\ell}}^{\mathrm{unip}}\mathbf{G}^{\mathrm{F}}$, applying $\langle\rho,$-$\rangle$ to 1, we have
$\sum_{\zeta^{\prime}}\langle\rho,d(\mathrm{IC}_{w}^{\mathrm{mix}}[q^{\frac{i}{2}}\zeta^{\prime}])\rangle=\sum_{\zeta^{\prime}}\langle\rho,d(\mathrm{IC}_{w}^{\mathrm{mix}})[q^{\frac{i}{2}}\zeta^{\prime}]\rangle.$ (2)
By Theorem 4.2.1, the LHS is
$\sum_{\zeta^{\prime}}\langle\rho,d(\mathrm{IC}_{w}^{\mathrm{mix}}[q^{\frac{i}{2}}\zeta^{\prime}])\rangle$ $=\sum_{\zeta^{\prime}}\langle d(\rho),\mathrm{IC}_{w}^{\mathrm{mix}}[q^{\frac{i}{2}}\zeta^{\prime}]\rangle$
$=\langle d(\rho),\mathrm{IC}_{w}^{\mathrm{mix}}[q^{\frac{i}{2}}\zeta_{d(\rho)}]\rangle$
$=\langle\rho,d(\mathrm{IC}_{w}^{\mathrm{mix}}[q^{\frac{i}{2}}\zeta_{d(\rho)}])\rangle.$
Similarly, the RHS evaluates to
$\sum_{\zeta^{\prime}}\langle\rho,d(\mathrm{IC}_{w}^{\mathrm{mix}})[q^{\frac{i}{2}}\zeta^{\prime}]\rangle=\langle\rho,d(\mathrm{IC}_{w}^{\mathrm{mix}})[q^{\frac{i}{2}}\zeta_{\rho}]\rangle.$
It thus follows from Lemma 4.2.2 that
$\langle\rho,d(\mathrm{IC}_{w}^{\mathrm{mix}}[q^{\frac{i}{2}}\zeta])\rangle=\langle\rho,d(\mathrm{IC}_{w}^{\mathrm{mix}})[q^{\frac{i}{2}}\zeta]\rangle$
for all $\zeta$ and the lemma follows. ∎
Corollary 4.2.4.
Let $A\in\mathcal{O}^{\mathrm{DL},\mathrm{unip},\mathrm{mix}}$, for all $\overline{\lambda}\in\overline{\mathbb{F}}_{\ell}^{\times}$, there is an equality in $K_{0}(\mathrm{Rep}_{\overline{\mathbb{Q}}_{\ell}}(\mathbf{G}^{\mathrm{F}}))$
$d(\mathrm{ch}(A)[\overline{\lambda}])=(d(\mathrm{ch}(A))[\overline{\lambda}]).$
Proof.
This is simply a matter of grouping terms and applying Lemma 4.2.3. ∎